OpenStudy (marcon):

Three functions are given below: f(x), g(x), and h(x). Explain how to find the axis of symmetry for each function, and rank the functions based on their axis of symmetry (from smallest to largest).

OpenStudy (marcon):

f(x) f(x) = –4(x – 8^)2 + 3 g(x) g(x) = 3x^2 + 12x + 15

OpenStudy (marcon):

h(x)

OpenStudy (marcon):

@jim_thompson5910 can you help me?

OpenStudy (marcon):

@cwrw238 @Loser66

jimthompson5910 (jim_thompson5910):

what do you have so far?

OpenStudy (marcon):

I don't know how to do the f(x) and g(x) but I think I get the h(x)

jimthompson5910 (jim_thompson5910):

what did you get for the h(x) function

OpenStudy (marcon):

Does it have to be an equation?

jimthompson5910 (jim_thompson5910):

yeah the axis of symmetry is an equation of the form x = h where h is some fixed number

OpenStudy (marcon):

I don't know how to do this...

jimthompson5910 (jim_thompson5910):

ok forget about the equation, what did you have for h(x) ?

jimthompson5910 (jim_thompson5910):

you wrote "I think I get the h(x)" what did you have in mind?

OpenStudy (marcon):

I was thinking it would be x=3 because that's where axis of symmetry crosses the parabola, but it has to be an equation which I don't know how to do.

jimthompson5910 (jim_thompson5910):

it is x = 3 and it's because this is the center line that divides the parabola in half (you can mirror one side over this line to get the other side)

jimthompson5910 (jim_thompson5910):

x = 3 is an equation anything that has an equal sign is an equation

OpenStudy (marcon):

Oh so h(x) would just be x=3?

jimthompson5910 (jim_thompson5910):

h(x) is some parabola we don't really care about the equation for h(x) but the equation for the axis of symmetry is x = 3 (the axis of symmetry of h(x))

OpenStudy (marcon):

hey I need to go somewhere, i'll tag you when I get back to finish the rest.

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (marcon):

I'm back @jim_thompson5910

jimthompson5910 (jim_thompson5910):

welcome back

OpenStudy (marcon):

So we got h(x). How do I do f(x) and g(x) ?

jimthompson5910 (jim_thompson5910):

f(x) is really f(x) = -4(x-8)^2 + 3 right?

OpenStudy (marcon):

yup

jimthompson5910 (jim_thompson5910):

f(x) = -4(x-8)^2 + 3 is in the form f(x) = a(x-h)^2 + k which is vertex form

jimthompson5910 (jim_thompson5910):

what is the value of h?

OpenStudy (marcon):

I just solve the equation to get h?

jimthompson5910 (jim_thompson5910):

you really don't need to solve you just need to compare and line up the numbers with the letters

OpenStudy (marcon):

I got (8,3) for the vertex

jimthompson5910 (jim_thompson5910):

the x coordinate of the vertex is what helps determine the axis of symmetry

OpenStudy (marcon):

Axis of symmetry is x = 8

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

for g(x), I recommend you graph it using a graphing tool like desmos https://www.desmos.com/calculator

OpenStudy (marcon):

For g(x) I got x=-2 as the axis of symmetry because the vertex is (-2,3) Is that correct?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (marcon):

So to rank them from smallest to largest it would be g(x) x=-2 , h(x) x=3, and f(x) x=8 right?

jimthompson5910 (jim_thompson5910):

I agree with that order

OpenStudy (marcon):

Thanks for your help!

jimthompson5910 (jim_thompson5910):

np